Complete the ordered pair ( ,4) so that it is the solution for x+2y=9.

Answer by KnightOwlTutor(293)

x+2y=9
We know the value of y which is 4.
Plug this value into the equation
x+2(4)=9
x+8=9
Subtract 8 from both sides
x=1
Therefore the answer is (1,4)